# Feature Representation Transferring to Lightweight Models via Perception Coherence

### v2

@@ Abstract
In this paper, we propose a method for transferring feature representation to lightweight student models from larger teacher models. We mathematically define a new notion called perception coherence. Based on this notion, we propose a loss function, which takes into account the dissimilarities between data points in feature space through their ranking. At a high level, by minimizing this loss function, the student model learns to mimic how the teacher model perceives inputs. More precisely, our method is motivated by the fact that the representational capacity of the student model is weaker than the teacher model. Hence, we aim to develop a new method allowing for a better relaxation. This means that, the student model does not need to preserve the absolute geometry of the teacher one, while preserving global coherence through dissimilarity ranking. Importantly, while rankings are defined only on finite sets, our notion of perception coherence extends them into a probabilistic form. This formulation depends on the input distribution and applies to general dissimilarity metrics. Our theoretical insights provide a probabilistic perspective on the process of feature representation transfer. Our experiments results show that our method outperforms or achieves on-par performance compared to strong baseline methods for representation transferring.
@@ 1 Introduction
The need of lightweight models and Knowledge distilling (KD). Deep learning models are at the forefront of performance in different domains and tasks, such as classification (He et al., 2016; Tan & Le, 2019; 2021; He et al., 2020) or object detection (Ren et al., 2015; Lin et al., 2017a; b; He et al., 2020). However, as the performance of neural networks increases, so does the cost of increasingly large models. In many applications where resources are limited (e.g. mobile devices) or one needs a fast execution, it is preferable to use lightweight models. Among the many approaches to obtain lightweight models, KD (Buciluǎ et al., 2006; Hinton et al., 2015; Wang & Yoon, 2021) stands out as a particularly interesting direction. A KD setup essentially consists of a teacher model and a student model. The student model (here the targeted lightweight model) is generally smaller than the teacher model, and therefore more efficient for the task in question. At a high level, the KD process consists in teaching the student model so that it mimics the way the teacher model understands the input. This is usually achieved by using a so-called transfer dataset. The student model is trained to behave like the teacher model on this set in some way . Many previous works have shown that KD improves the generalization capacity of the student model (Hinton et al., 2015; Chen et al., 2017; Zhu et al., 2018; Beyer et al., 2022; Huang et al., 2022). Different up-to-date KD approaches will be discussed in Section 2.
@@
Feature representation transferring. Classical KD techniques generally focus on matching the outputs of student and teacher models, typically for classification tasks (Hinton et al., 2015). However, this approach has limitations, such as requiring the same number of classes between the teacher and student models. To address these issues, previous works (Passalis & Tefas, 2018a; Passalis et al., 2020b; Tian et al., 2020; Park et al., 2019) propose to directly transfer the feature space representations through learning. This enables the student model to better capture the input-input relations represented in the feature space of the teacher model, by preserving its geometry and distribution (illustrated in Fig. 1). This leads to more effective knowledge transfer. Furthermore, note that this approach is an unsupervised transferring method, as it only uses a transfer set without labels. As a consequence, the learned features can be used for various downstream tasks. Because of these appealing properties, our method aligns with this last approach.
@@
  • We conduct experiments on retrieval and classification tasks to compare our method with strong baseline methods. The results show that our method outperforms or achieves competitive performance with strong baseline methods on lightweight representation transferring settings. Our code is available at anonymized-code.
@@ 2 Related Work
KD based on soft label produced by teacher model. This approach has been studied since the early days of neural networks (Buciluǎ et al., 2006). Therein, the authors show that knowledge can be transferred from a large ensemble of models to a single small model. A standard KD method consists in using output of classification models (Hinton et al., 2015; Tang et al., 2016; Huang et al., 2022), with a temperature, leading to smoother soft labels. Based on the same spirit, (Chan et al., 2015) proposes to match the output distributions using Kullback-Leibler divergence. Soft label was also used in the context of domain adaption (Tzeng et al., 2015). While all these methods show the benefits of soft labels, the main limit of this approach is the need of having the same number of classes between models.
@@
KD based on distance matching. In Romero et al. (2014), the authors propose to minimize the distance between student and teacher’s features to guide the student model during training. However, in the case of dimension mismatch between the models, one need to pass through a linear transformation, which leads to substantial information loss. In Yim et al. (2017), the authors propose to minimize distance of the so-call FSP matrix between teacher and student models (capturing inter-layer relations). More advanced methods consist in minimizing the distance between the derivatives w.r.t. to the features (Czarnecki et al., 2017) or between the attention map of student and teacher (Zagoruyko & Komodakis, 2016). All these approaches require the dimension match, and need some form of interpolation in the case of dimension mismatch. Overcoming the drawback of dimension mismatch, Passalis & Tefas (2018b) proposes to match the the pair-wise similarity matrix of the training points (in the transfer set) between the student and teacher models. The method proposed in Yu et al. (2019) is to match the distance between pairs of samples. However, this brute force approach is not likely to work as the student model is much smaller than the teacher one (as discussed in Passalis et al. (2020b)).
@@
Feature representation transfer. Some previous works treat the representation transfer as the problem of distribution matching, based on Maximum Mean Discrepancy (Huang & Wang, 2017) or adversarial training (Belagiannis et al., 2018). The latter method is inspired by Generative Adversarial Networks (GANs) Goodfellow et al. (2014). One major drawback of distribution matching is the need of having the same dimension between the two feature spaces. Probabilistic frameworks based on mutual information are proposed in Ahn et al. (2019); Tian et al. (2020). However, the method of Ahn et al. (2019) requires a Gaussian assumption (this is not necessarily satisfied in practice) while the method of Tian et al. (2020) requires the input features to have the same dimension (hence requiring a linear transformation, which may reduce the useful information). Overcoming the requirement of dimension match, Park et al. (2019) proposes to transfer the input relations using distance-wise and angle-wise distillation loss. However, the structural information is quite limited, as it only considers a single triple of points at a time. Overcoming this drawback, Passalis & Tefas (2018a) proposes a method of probabilistic knowledge transfer (PKT). The high level idea of this last method is to transfer the local geometry learned by the teacher model based on kernel techniques. Based on this work, Passalis et al. (2020a; b) basically share the same idea, but they apply the technique for multiples intermediate layers with an auxiliary model. The benefits of using auxiliary models are discussed in Mirzadeh et al. (2020). As in Passalis & Tefas (2018a); Passalis et al. (2020a; b), our method takes into account the relation between the inputs and allows representation transfer between space having different dimensions. However, in Passalis & Tefas (2018a); Passalis et al. (2020a; b), a common kernel is used for all points (e.g. a Gaussian one with a fixed bandwidth), which cannot adapt to local density variation. This contrasts with our method, where we only rank the dissimilarities but do not use the dissimilarity magnitude. Thus, by using the dissimilarity rankings (or equivalently cumulative function of dissimilarities), our method captures local probabilistic closeness — a property that raw distance magnitudes fail to reflect in the presence of density variations.
@@ 3.2 Practical implementation with a differentiable ranking operation
To maximize the perception coherence level (Eq. (2)), we need to somehow minimize the discrepancy between $F_{1}$ and $F_{2}$. Thus, empirically, we can minimize
@@ 4.1 Expected Convergence of the Mini-Batch Estimator
Mini-batch approximation.
In practice, computing $\mathrm{DC}$ exactly is infeasible, as it requires access to the full distribution (or the entire dataset). Instead, we approximate it using i.i.d. samples drawn in mini-batches. The corresponding empirical estimator is defined as follows.
@@
Convergence of the mini-batch estimator.
We now establish that $\widehat{\mathrm{DC}}_{B}$ converges to the true $\mathrm{DC}$ at the standard Monte Carlo rate.
@@
Discussion.
Theorem 4.1 ensures that the estimation error decays proportionally to $1/\sqrt{B}$. In other words, enlarging the mini-batch size systematically reduces the variance of the estimator, thereby improving the accuracy of the approximation to $\mathrm{DC}$. This theoretical guarantee is strongly supported by our experiments, that will be shown in Section 5.4 (Fig. 6 and Table 4). More precisely, we shall show that the empirical estimation error exhibits the predicted scaling behavior with respect to $B$.
@@ 4.4 Stability of perception coherence around a local region
Consider a non-empty set $\mathcal{A}\subseteq\mathcal{X}$ such that $\phi_{f_{1},f_{2}}(x)\geq\alpha,\;\forall x\in\mathcal{A}$. For $0<\delta\leq 1$, let $\mathcal{A}_{\delta}=\{x\in\mathcal{X}:\min_{x^{\prime}\in\mathcal{A}}F(x^{\prime},x)\leq\delta\}$. Then, for any $0<\varepsilon\leq\alpha$, we have
@@ 5 Experiments
In this section, we first conduct experiments on $2D$ and $3D$ data for proof-of-concept (Section 5.1). Then, in Section 5.2, we conduct experiments on lightweight settings to evaluate the quality of the learned representation through retrieval tasks. Finally, we demonstrate how our method can boost the performance of student models in classification task in Section 5.3. We shall also conduct ablation study in Sections 5.4 and 5.5. Without further mention, we use cosine dissimilarity as the dissimilarity metric on neural networks (defined in Appendix E). For brevity, the details for all experiments can also be found in Appendix E.
@@ 5.2 Learning Metric Evaluation on Neural Networks
In this section, we perform experiments on CIFAR10 Krizhevsky et al. (2009) and the subset of the first 30 classes of CUB-200 Wah et al. (2011) with very small student models. Following the previous works (e.g. Passalis & Tefas (2018a); Passalis et al. (2020b; a)), we evaluate the quality of the learned representation using the content-based image retrieval setup. We used the official code released with the work Passalis et al. (2020a) to evaluate different metrics, including the interpolated mean Average Precision (mAP) (at the standard 11-recall points) and the top-$k$ precision (Christopher et al., 2008). For a fair comparison, we use the same teacher and student model architecture as in Passalis et al. (2020a). The student models are very small, composed of only 3 convolutional layers and a fully connected layer. All details are in Appendix E.2. To conduct the experiments, for each dataset CIFAR10 and CUB-200, we first train the teacher model on the training set. Then, the transfer set is only the training set without label. Then, the retrieval setup is used to evaluate the quality of the learned representation. The database contains the training set and the test set is used for evaluating the performance of each model. If the student is trained to learn useful features, then the test examples should be close to the training examples (database) of the same class. Hence, a higher retrieval score indicates a better quality of the learned representation. The results for CIFAR10 and CUB-200 are shown in Tables 1 and 2, respectively. Our method is also compared with classical methods, including the standard knowledge distilling (KD) Hinton et al. (2015), FitNet Romero et al. (2014), metric knowledge transfer (MKT) Yu et al. (2019), probabilistic knowledge transfer (PKT) Passalis & Tefas (2018a) and Heterogeneous Knowledge Distillation (HKD) Passalis et al. (2020a). The results displayed for all these methods in Tables 1 and 2 are extracted from Passalis et al. (2020a). Therein, one used the same teacher and student models. For FitNet, MKT, PKT, there are two versions: the first one applies only on a single penultimate layer, while the second one applies on multiple layers (denoted by -H).
@@
From Tables 1 and 2, we observe that MKT method provides the worst results. This method is composed of 2 main ingredients: (1) minimizing the distance between the student and teacher features and (2) enforcing the student model to result in similar distances between the data points as the teacher model (in the feature space). Hence, in our setting of very small student models, this brute force approach provides the worst results. Our method also outperforms by a large gap the standard approaches of KD (using the soft label) or FitNet (using the distance between the student and teacher features). More advanced methods such as PKT (based on a kernel method) performs better, but it is still outperformed by our method. Besides, HKD is a more advanced version of PKT, where one uses a particular mechanism to transfer the features. However, our method still provides better results. Notice that our method is applied only on the penultimate layer, but outperforms the other methods using multiple layers. Especially on CUB-200, our method outperforms the others by a wide margin. This proves the effectiveness of our method for feature representation transferring.
@@
In the above section, we show how our method helps the student model to produce useful features. In this section, we demonstrate that by transferring the feature representation, our method boosts significantly the performance of the student model in the context of the classification task on CIFAR100 Krizhevsky (2009). While different methods have been proposed and tailored specifically for the classification, we focus mainly on state-of-the-art feature representation transferring methods. Notice that different methods (such as FitNet Romero et al. (2014)) apply the transferring on different intermediate layers. However, we propose to apply our method only on the penultimate layer and on the logit one (before applying the softmax function). We also fix all the hyperparameters for training, without further tuning to show the simplicity of our method. All details can be found in Appendix E.3. While different methods have been proposed to effectively transfer knowledge between model of the same architecture type (e.g. ResNet50 $\to$ ResNet18), we intentionally choose pairs of models having different architecture types. More precisely, we use 3 pairs of teacher and student models: ResNet-50 $\to$ MobileNetV2 and ResNet-32x4 $\to$ ShuffleNetV1 / ShuffleNetV2. We follow the same setup and use the same pretrained teacher models officially released with the work Tian et al. (2020) for fair comparisons.
Table 3: Test classification accuracy ($\%$) on CIFAR100 using different knowledge distillation methods. Results of other methods are extracted from Huang et al. (2022), where one used the same setups.
@@
From Table 3, we see that our method boosts significantly the performance of the student models (when not using any KD technique). It also provides much better results than standard techniques such as KD or FitNet. This proves its effectiveness for extracting knowledge from the teacher model. Moreover, it outperforms all the other methods on ResNet-32x4/ShuffleNetV1 and ResNet-32x4 /ShuffleNetV2 benchmarks. On the experiment ResNet-50/ MobileNetV2, our method is slightly outperformed by the method CRD, but outperforms all the other methods. Besides, our method systematically outperforms the PKT method based on kernel techniques (see Section 2). This reinforces our intuition that our method is more effective than applying the same kernel for all points, as discussed in Section 2. Other methods based on mutual information such as Variational Information Distillation (VID) is also outperformed by ours. The method of Contrastive Representation Distillation (CRD) is also based on an information-theoretic framework. Our method provides better or on-par performance compared to CRD. All these remarks prove the effectiveness of our method for KD.

### v3

@@ Abstract
In this paper, we propose a method for transferring feature representation to lightweight student models from larger teacher models. We mathematically define a new notion called perception coherence. Based on this notion, we propose a loss function, which takes into account the dissimilarities between data points in feature space through their ranking. At a high level, by minimizing this loss function, the student model learns to mimic how the teacher model perceives inputs. More precisely, our method is motivated by the fact that the representational capacity of the student model is weaker than the teacher model. Hence, we aim to develop a conceptually new method allowing for a better relaxation. This means that, the student model does not need to preserve the absolute geometry of the teacher one, while preserving global coherence through dissimilarity ranking. Importantly, while rankings are defined only on finite sets, our notion of perception coherence extends them into a probabilistic form. This formulation depends on the input distribution and applies to general dissimilarity metrics. Our theoretical insights provide a probabilistic perspective on the process of feature representation transfer. Our experimental results show that our method outperforms or achieves on-par performance with strong baseline methods for representation transfer, particularly class-unaware ones.
@@ 1 Introduction
The need of lightweight models and Knowledge distilling (KD). Deep learning models are at the forefront of performance in different domains and tasks, such as classification (He et al., 2016; Tan and Le, 2019; 2021; He et al., 2020) or object detection (Ren et al., 2015; Lin et al., 2017a; b; He et al., 2020). However, as the performance of neural networks increases, so does the cost of increasingly large models. In many applications where resources are limited (e.g. mobile devices) or one needs a fast execution, it is preferable to use lightweight models. Among the many approaches to obtain lightweight models, KD (buciluǎ2006model; Hinton et al., 2015; Wang and Yoon, 2021) stands out as a particularly interesting direction. A KD setup essentially consists of a teacher model and a student model. The student model (here the targeted lightweight model) is generally smaller than the teacher model, and therefore more efficient for the task in question. At a high level, the KD process consists in teaching the student model so that it mimics the way the teacher model understands the input. This is usually achieved by using a so-called transfer dataset. The student model is trained to behave like the teacher model on this set in some way . Many previous works have shown that KD improves the generalization capacity of the student model (Hinton et al., 2015; Chen et al., 2017; Zhu et al., 2018; Beyer et al., 2022; Huang et al., 2022). Different up-to-date KD approaches will be discussed in Section 2.
@@
Feature representation transferring. Classical KD techniques generally focus on matching the outputs of student and teacher models, typically for classification tasks (Hinton et al., 2015). However, this approach has limitations, such as requiring the same number of classes between the teacher and student models. To address these issues, previous works (Passalis and Tefas, 2018a; Passalis et al., 2020b; Tian et al., 2020; Park et al., 2019) propose to directly transfer the feature space representations through learning. This enables the student model to better capture the input-input relations represented in the feature space of the teacher model, by preserving its geometry and distribution (illustrated in Fig. 1). This leads to more effective knowledge transfer. Furthermore, note that this approach is an unsupervised transferring method, as it only uses a transfer set without labels. As a consequence, the learned features can be used for various downstream tasks. Because of these appealing properties, our method aligns with this last approach.
@@
  • We conduct experiments on retrieval and classification tasks to compare our method with strong baseline methods. The results show that our method outperforms or achieves competitive performance with strong baseline methods on lightweight representation transferring settings. Our code is available at perception-coherence-code.
@@ 2 Related Work
KD based on soft label produced by teacher model. This approach has been studied since the early days of neural networks (buciluǎ2006model). Therein, the authors show that knowledge can be transferred from a large ensemble of models to a single small model. A standard KD method consists in using output of classification models (Hinton et al., 2015; Tang et al., 2016; Huang et al., 2022), with a temperature, leading to smoother soft labels. Based on the same spirit, (Chan et al., 2015) proposes to match the output distributions using Kullback-Leibler divergence. Soft label was also used in the context of domain adaption (Tzeng et al., 2015). While all these methods show the benefits of soft labels, the main limit of this approach is the need of having the same number of classes between models.
@@
KD based on distance matching. In Romero et al. (2014), the authors propose to minimize the distance between student and teacher’s features to guide the student model during training. However, in the case of dimension mismatch between the models, one need to pass through a linear transformation, which leads to substantial information loss. In Yim et al. (2017), the authors propose to minimize distance of the so-call FSP matrix between teacher and student models (capturing inter-layer relations). More advanced methods consist in minimizing the distance between the derivatives w.r.t. to the features (Czarnecki et al., 2017) or between the attention map of student and teacher (Zagoruyko and Komodakis, 2016). All these approaches require the dimension match, and need some form of interpolation in the case of dimension mismatch. Overcoming the drawback of dimension mismatch, Passalis and Tefas (2018b) proposes to match the the pair-wise similarity matrix of the training points (in the transfer set) between the student and teacher models. The method proposed in Yu et al. (2019) is to match the distance between pairs of samples. However, this brute force approach is not likely to work as the student model is much smaller than the teacher one (as discussed in Passalis et al. (2020b)).
@@
Feature representation transfer. Some previous works treat the representation transfer as the problem of distribution matching, based on Maximum Mean Discrepancy (Huang and Wang, 2017) or adversarial training (Belagiannis et al., 2018). The latter method is inspired by Generative Adversarial Networks (GANs) Goodfellow et al. (2014). One major drawback of distribution matching is the need of having the same dimension between the two feature spaces. Probabilistic frameworks based on mutual information are proposed in Ahn et al. (2019); Tian et al. (2020). However, the method of Ahn et al. (2019) requires a Gaussian assumption (this is not necessarily satisfied in practice) while the method of Tian et al. (2020) requires the input features to have the same dimension (hence requiring a linear transformation, which may reduce the useful information). Overcoming the requirement of dimension match, Park et al. (2019) proposes to transfer the input relations using distance-wise and angle-wise distillation loss. However, the structural information is quite limited, as it only considers a single triple of points at a time. Overcoming this drawback, Passalis and Tefas (2018a) proposes a method of probabilistic knowledge transfer (PKT). The high level idea of this last method is to transfer the local geometry learned by the teacher model based on kernel techniques. Based on this work, Passalis et al. (2020a; b) basically share the same idea, but they apply the technique for multiples intermediate layers with an auxiliary model. The benefits of using auxiliary models are discussed in Mirzadeh et al. (2020). As in Passalis and Tefas (2018a); Passalis et al. (2020a; b), our method takes into account the relation between the inputs and allows representation transfer between space having different dimensions. However, in Passalis and Tefas (2018a); Passalis et al. (2020a; b), a common kernel is used for all points (e.g. a Gaussian one with a fixed bandwidth), which cannot adapt to local density variation. This contrasts with our method, where we only rank the dissimilarities but do not use the dissimilarity magnitude. Thus, by using the dissimilarity rankings (or equivalently cumulative function of dissimilarities), our method captures local probabilistic closeness — a property that raw distance magnitudes fail to reflect in the presence of density variations.
Class-awareness. Knowledge distillation is a broad research domain, known for leveraging the “knowledge” of a teacher model to improve a student model. However, most state-of-the-art methods adopt a class-aware approach. This means that these methods rely on class information in some form—most notably, the number of classes. The standard approach involves matching the output logits (Hinton et al., 2015; Huang et al., 2022). More recent class-aware methods reuse the teacher’s classifier (Chen et al., 2022) and then retrain the student encoder. Other methods employ the teacher’s classification head to implicitly align the student’s intermediate representations with those of the teacher (Wang et al., 2024). While such methods demonstrate strong performance on certain benchmarks, they are often tailored to specific tasks. In contrast, our method is conceptually designed for general-purpose representation transfer, applicable to any feature space equipped with a pairwise dissimilarity measure. This class-unaware formulation conceptually enables broader applicability, e.g., regression models or even hand-crafted features.
@@ 3.2 Practical implementation with a differentiable ranking operation
To maximize the perception coherence level, we need to somehow minimize the discrepancy between $F_{1}$ and $F_{2}$. Thus, empirically, we can minimize
@@ 4.1 Expected Convergence of the Mini-Batch Estimator
Mini-batch approximation. In practice, computing $\mathrm{DC}$ exactly is infeasible, as it requires access to the full distribution (or the entire dataset). Instead, we approximate it using i.i.d. samples drawn in mini-batches. The corresponding empirical estimator is defined as follows.
@@
Convergence of the mini-batch estimator. We now establish that $\widehat{\mathrm{DC}}_{B}$ converges to the true $\mathrm{DC}$ at the standard Monte Carlo rate.
@@
Discussion. Theorem 4.1 ensures that the estimation error decays proportionally to $1/\sqrt{B}$. In other words, enlarging the mini-batch size systematically reduces the variance of the estimator, thereby improving the accuracy of the approximation to $\mathrm{DC}$. This theoretical guarantee is strongly supported by our experiments, that will be shown in Section 5.5 (Fig. 7 and Table 5). More precisely, we shall show that the empirical estimation error exhibits the predicted scaling behavior with respect to $B$.
@@ 4.4 Stability of perception coherence around a local region
Consider a non-empty set $\mathcal{A}\subseteq\mathcal{X}$ (with $\mathbb{P}_{X}\left(X\in\mathcal{A}\right)>0$) such that $\phi_{f_{1},f_{2}}(x)\geq\alpha,\;\forall x\in\mathcal{A}$. For $0<\delta\leq 1$, let $\mathcal{A}_{\delta}=\{x\in\mathcal{X}:\min_{x^{\prime}\in\mathcal{A}}F(x^{\prime},x)\leq\delta\}$. Then, for any $0<\varepsilon\leq\alpha$, we have
@@ 5 Experiments
In this section, we first conduct experiments on $2D$ and $3D$ data for proof-of-concept (Section 5.1). Next, in Section 5.2, we present a stylized empirical study showing the correlation between the perception coherence level during transfer and the performance of downstream classification task. Then, in Section 5.3, we conduct experiments on lightweight settings to evaluate the quality of the learned representation through retrieval tasks. Besides, we demonstrate how our method can boost the performance of student models in classification task in Section 5.4. Note that in all experiments conducted in Sections 5.3 and 5.4, the student models are smaller than the teacher models [^2] See Appendix F for detailed computational cost comparison of different pairs of teacher-student models., to meet practical requirements. Finally, we shall also conduct ablation study in Sections 5.5 and 5.6. Without further mention, we use cosine dissimilarity as the dissimilarity metric on neural networks (defined in Appendix E). For brevity, the details for all experiments can also be found in Appendix E.
@@ 5.1 Proof-of-concept with experiments on $\mathbf{2D}$ and $\mathbf{3D}$ toy examples
5.2 A Stylized Empirical Study of Perception Coherence and Downstream Classification
To show the correlation between perception coherence and downstream task performance, we study its relationship with classification performance in a controlled setting. We consider a two-moon binary classification dataset with 800 samples, randomly split into 400 training and 400 test points. For teacher model [^3] For brevity, please refer to Appendix E.2 for model details, including training details., a neural network with two hidden layers followed by a softmax classifier is trained in a fully supervised manner, using the labeled training set. After convergence, the teacher achieves perfect accuracy on both training and test sets.
Student model and unsupervised transfer. For simplicity, the student network shares the same architecture as the teacher model. Using our proposed unsupervised transfer method, we distill the representation from the teacher’s penultimate layer into the student, relying only on unlabeled training data. Importantly, the final softmax layer of the student is excluded from this transfer process. During training, we save student checkpoints at ten different epochs, denoted by $\{S_{k}\}_{k=1}^{10}$.
Perception coherence. For each checkpoint $S_{k}$, we compute the perception coherence coefficient on the training set, yielding values $\{\mathrm{PC}_{k}\}_{k=1}^{10}$.
Downstream classification. To evaluate the usefulness of the transferred representations, we freeze all feature layers of each student model in $\{S_{k}\}_{k=1}^{10}$ and train only the final softmax layer in a supervised fashion on the labeled training set. This protocol isolates the quality of the learned representation: higher accuracy directly reflects task-relevant features. After training, by testing the trained models on the test set, we obtain the test accuracies $\{\mathrm{Acc}_{k}\}_{k=1}^{10}$ corresponding to different checkpoints $\{S_{k}\}_{k=1}^{10}$.
Table 1 and Fig. 6 show the relation between the perception coherence values and corresponding downstream classification accuracies (i.e., $\{\mathrm{PC}_{k}\}_{k=1}^{10}$ v.s. $\{\mathrm{Acc}_{k}\}_{k=1}^{10}$). We observe a strong positive relationship between perception coherence and downstream performance: student models with higher PC consistently achieve better classification accuracy. Quantitatively, the Pearson correlation coefficient between training PC and test accuracy is $0.920$. These results establish a direct task-level connection between perception coherence and classification error. Even in this stylized setting, improved coherence of the learned representation leads to higher linear separability and better downstream accuracy. This supports the interpretation of perception coherence as a meaningful proxy for representation transferring quality (without the need for label).
Table 1: Training perception coherence versus downstream test accuracy for student models at different training epochs.
Training perception coherence level 0.841 0.877 0.889 0.901 0.911 0.921 0.931 0.940 0.949 0.956
Test accuracy (%) 82.00 83.50 83.25 85.75 86.75 88.00 86.75 89.25 87.00 90.00
Figure 6: Training perception coherence versus downstream test accuracy for student models at different training epochs.
@@ 5.3 Learning Metric Evaluation on Neural Networks
In this section, we perform experiments on CIFAR10 (Krizhevsky et al., 2009) and the subset of the first 30 classes of CUB-200 (Wah et al., 2011) with very small student models. Following the previous works (e.g. (Passalis and Tefas, 2018a; Passalis et al., 2020b; a)), we evaluate the quality of the learned representation using the content-based image retrieval setup. We used the official code released with the work of Passalis et al. (2020a) to evaluate different metrics, including the interpolated mean Average Precision (mAP) (at the standard 11-recall points) and the top-$k$ precision (Christopher et al., 2008). For a fair comparison, we use the same teacher and student model architecture as in Passalis et al. (2020a). The student models are very small, composed of only 3 convolutional layers and a fully connected layer. All details are in Appendix E.3. To conduct the experiments, for each dataset CIFAR10 and CUB-200, we first train the teacher model on the training set. Then, the transfer set is only the training set without label. Then, the retrieval setup is used to evaluate the quality of the learned representation. The database contains the training set and the test set is used for evaluating the performance of each model. If the student is trained to learn useful features, then the test examples should be close to the training examples (database) of the same class. Hence, a higher retrieval score indicates a better quality of the learned representation. The results for CIFAR10 and CUB-200 are shown in Tables 2 and 3, respectively. Our method is also compared with classical methods, including the standard knowledge distilling (KD) (Hinton et al., 2015), FitNet (Romero et al., 2014), metric knowledge transfer (MKT) (Yu et al., 2019), probabilistic knowledge transfer (PKT) (Passalis and Tefas, 2018a) and Heterogeneous Knowledge Distillation (HKD) (Passalis et al., 2020a). The results displayed for all these methods in Tables 2 and 3 are extracted from Passalis et al. (2020a). Therein, one used the same teacher and student models. For FitNet, MKT, PKT, there are two versions: the first one applies only on a single penultimate layer, while the second one applies on multiple layers (denoted by -H).
@@
From Tables 2 and 3, we observe that MKT method provides the worst results. This method is composed of 2 main ingredients: (1) minimizing the distance between the student and teacher features and (2) enforcing the student model to result in similar distances between the data points as the teacher model (in the feature space). Hence, in our setting of very small student models, this brute force approach provides the worst results. Our method also outperforms by a large gap the standard approaches of KD (using the soft label) or FitNet (using the distance between the student and teacher features). More advanced methods such as PKT (based on a kernel method) performs better, but it is still outperformed by our method. Besides, HKD is a more advanced version of PKT, where one uses a particular mechanism to transfer the features. However, our method still provides better results. Notice that our method is applied only on the penultimate layer, but outperforms the other methods using multiple layers. Especially on CUB-200, our method outperforms the others by a wide margin. This proves the effectiveness of our method for feature representation transferring.
@@
In the above section, we show how our method helps the student model to produce useful features. In this section, we demonstrate that by transferring the feature representation, our method boosts significantly the performance of the student model in the context of the classification task on CIFAR100 (Krizhevsky, 2009). While different methods have been proposed and tailored specifically for the classification, we focus mainly on state-of-the-art feature representation transferring methods. Notice that different methods (such as FitNet (Romero et al., 2014)) apply the transferring on different intermediate layers. However, we propose to apply our method only on the penultimate layer and on the logit one (before applying the softmax function). We also fix all the hyperparameters for training, without further tuning to show the simplicity of our method. The most recent VRM method (Zhang et al., 2025) introduces virtual relation matching and applies RandAug (Random Augmentation). For a fair comparison, we add RandAug to the inputs using the official code associated with this last work. Then, we apply our method under the same setting. All details can be found in Appendix E.4. While different methods have been proposed to effectively transfer knowledge between model of the same architecture type (e.g. ResNet50 $\to$ ResNet18), we intentionally choose pairs of models having different architecture types. More precisely, we use 3 pairs of teacher and student models: ResNet-50 $\to$ MobileNetV2 and ResNet-32x4 $\to$ ShuffleNetV1 / ShuffleNetV2. We follow the same setup and use the same pretrained teacher models officially released with the work of Tian et al. (2020) for fair comparisons.
Table 4: Test classification accuracy ($\%$) on CIFAR100 using different knowledge distillation methods. Results of ReviewKD (Chen et al., 2021), TTM (Zheng and YANG, 2024) and VRM (Zhang et al., 2025) are extracted from the corresponding original papers, other methods are extracted from Huang et al. (2022), where one used the same setups.
@@
TTM (Zheng and YANG, 2024) 69.59 74.37 76.55
VRM (with RandAug) (Zhang et al., 2025) 72.30 78.28 79.34
[…]
From Table 4, we see that our method boosts significantly the performance of the student models (when not using any KD technique). It also provides much better results than standard techniques such as KD or FitNet. This proves its effectiveness for extracting knowledge from the teacher model. Besides, our method systematically outperforms the PKT method based on kernel techniques (see Section 2). This reinforces our intuition that our method is more effective than applying the same kernel for all points, as discussed in Section 2. Other methods based on mutual information such as Variational Information Distillation (VID) is also outperformed by ours. The method of Contrastive Representation Distillation (CRD) is also based on an information-theoretic framework. Our method outperforms this last one. Moreover, it outperforms all the other methods on ResNet-32x4/ShuffleNetV1 benchmark. We also observe that our method achieves on-par performance compared to VRM, even though VRM includes additional components such as inter-class relation modeling and edge pruning strategies. Our approach intentionally uses a minimal setting, yet still matches VRM performance. All these remarks prove the effectiveness of our method for KD.